# Zero-Shot Learning of Causal Models

### v2

@@ 5.2 Benchmark of Cond-FiP
An advantage of zero-shot inference methods is their generalization capabilities when the dataset size becomes smaller. As the test dataset size decreases, the performances of baselines can be severely affected since they require training from scratch on these datasets. In contrast, zero-shot inference methods are less impacted as their parameters remains unchanged at inference, and the inductive bias learned from training enables them to generalize even with a smaller input context. Appendix D provides detailed results for this case (illustration for specific setting in Figure 3), we find that Cond-FiP exhibits superior generalization as compared to the baselines.
Ablation Study.
We conduct ablations on the encoder and the decoder to better understand the effect of the training data on the generalization performances. We find that Cond-FiP with encoder trained using only linear/rff data is still competitive to the case of training on combined linear and rff data. However, for decoder, training on combined data is better than training with only linear/rff data. Please refer to Section E.1 and E.2 for more details.
We also test the sensitivity of Cond-FiP to distribution shifts in Section E.3, comparing its performance across scenarios as the severity of the distribution shift is increased. We consider two cases, where we control the severity in distribution shift by controlling the causal mechanisms or the noise variables. We find that Cond-FiP is robust to the varying levels of distribution shifts in causal mechanisms, as its performance doesn’t change much. However, Cond-FiP is quite sensitive to shifts varying level of distribution shifts in noise variables, with its performance degrading with increasing magnitude of shift in the noise variables.
6 Conclusion
In this work, we demonstrate that a single model can be trained to infer Structural Causal Models (SCMs) in a zero-shot manner through amortized training. Our proposed method, Cond-FiP, not only generalizes effectively to novel SCMs at test time but also remains robust across varying SCM distributions. To the best of our knowledge, this is the first approach to establish the feasibility of learning causal generative models in a foundational manner.
While Cond-FiP generalizes to new instance problems with larger graphs, it is still unable to improve its performances when given larger context sizes at inference (Appendix G). To further enhance the generalization capabilities, we would need to scale both the model and the training data, allowing the model to encounter more complex and diverse contexts. Future work will focus on scaling Cond-FiP to larger problem instances and application to real-world scenarios.
Acknowledgements
We thanks the members of the Machine Intelligence team at Microsoft Research for helpful discussions. We also thank Sarthak Mittal for their suggestion to benchmark Cond-FiP in the scarce data regime. Part of the experiments were enabled by the Digital Research Alliance of Canada (https://alliancecan.ca/en) and Mila cluster (https://docs.mila.quebec/index.html).
Appendix A Additional Details on Cond-FiP
A.1 Details on Encoder Training
We provide further details on training the encoder and show how recovering the noise is equivalent to learn the inverse causal generative process. Recall that an SCM is an implicit generative model that, given a noise sample N, generates the corresponding observation according to the following fixed-point equation in X
$\textbf{X}=F(\textbf{X},\textbf{N})$
$\textbf{X}=F_{\text{gen}}(\textbf{N})$
$\textbf{N}=F_{\text{gen}}^{-1}(\textbf{X})$
$F_{\text{gen}}^{-1}(z)=z-f(z)$
Additive Noise Model Assumption.
Our method relies on the ANM assumption only for the training the encoder. This is because we require the encoder to predict the noise from data in order to obtain embeddings, and under the ANM assumption, the mapping from data to noise can be easily expressed as $x\to x-F(x)$ where $F$ is the generative functional mechanism of the generative ANM. However, if we were to consider general SCMs, i.e. of the form $X=F(X,N)$, we would need access to the mapping $x\to F^{-1}(x,\cdot)(x)$ (assuming this function is invertible), which for general functions is not tractable. An interesting future work would be to consider a more general dataset encoding (using self-supervised techniques), but we believe this is out of the scope of this work.
A.2 Inference with Cond-FiP
Cond-FiP Model.
Once Cond-FiP is trained, we have access to two trained models: (1) an encoder that given a dataset $D_{\textbf{X}}$ and the causal graph associated $\mathcal{G}$ produced an embedding $\mu(D_{\textbf{X}},\mathcal{G})$ as defined l.277., and (2) a decoder that given an embedding $\mu$ and a causal graph $\mathcal{G}$, produces a function $z\in\mathbb{R}^{d}\to\mathcal{T}(z,\mu,\mathcal{G})\in\mathbb{R}^{d}$.
Cond-FiP simply consists of the composition of the two models, that is given a dataset $D_{\textbf{X}}$ and the causal graph associated $\mathcal{G}$, produce a function $z\in\mathbb{R}^{d}\to\mathcal{T}(z,\mu(D_{\textbf{X}},\mathcal{G}),\mathcal{G})\in\mathbb{R}^{d}$, which for simplicity we denote $z\in\mathbb{R}^{d}\to\mathcal{T}(z,D_{\textbf{X}},\mathcal{G})\in\mathbb{R}^{d}$. Having clarified this, we can now proceed to more detailed explanations of the inference process using Cond-FiP.
Sample Generation.
$\textbf{x}=\mathcal{T}(\textbf{x},D_{\bm{X}},\mathcal{G})+\textbf{n}$
$\bm{z}_{\ell}=\mathcal{T}(\bm{z}_{\ell-1},D_{\bm{X}},\mathcal{G})+\textbf{n}$
After $d$ iterations we obtain the following,
$\bm{z}_{d}=\mathcal{T}(\bm{z}_{d},D_{\bm{X}},\mathcal{G})+\textbf{n}$
Interventional Prediction.
Appendix B Experiments on Real World Benchmark
We use the real world flow cytometry dataset (Sachs et al., 2005) to benchmark Cond-FiP againts the baselines. This dataset contains $n\simeq 800$ observational samples expressed in a $d=11$ dimensional space, and the reference (true) causal graph. We sample a train dataset $D^{\textnormal{train}}_{\bm{X}}\in{\mathbb{R}}^{n_{\text{train}}\times d}$ and test dataset $D^{\textnormal{test}}_{\bm{X}}\in{\mathbb{R}}^{n_{\text{test}}\times d}$ of size $n_{\text{train}}=n_{\text{test}}=400$ each, where the train dataset is to used to train the baselines and obtain dataset embedding for Cond-FiP.
Since we don’t have access to the true functional relationships, we cannot compute RMSE for noise prediction or sample generation like we did in our experiments with synthetic benchmarks. Instead for each method, we obtain the noise predictions $\widehat{D^{\textnormal{train}}_{\bm{N}}}$ on the train split, and use it to fit a gaussian distribution for each component (node). Then we use the learned gaussian distribution to sample new noise variables, $\widehat{D^{\text{sample}}_{\bm{N}}}$, which are mapped to the observations as per the causal mechanisms learned by each method, $\widehat{D^{\text{sample}}_{\bm{X}}}$. Finally, we compute the maximum mean discrepancy (MMD) distance between $\widehat{D^{\text{sample}}_{\bm{X}}}$ and $D^{\textnormal{test}}_{\bm{X}}$ as metric to determine whether the method has captured the true causal relationships. For consistency, we also evaluate the reconstruction performances of the models by using directly the inferred noise $\widehat{D^{\textnormal{train}}_{\bm{N}}}$ from the models, and the compute MMD between the reconstructed data and the test data.
Table 4 presents our results, where for reference we also report the MMD distance between the true train and test split, which should be very small since both the datasets are sampled from the same distribution. We find that Cond-FiP is competitive with the baselines that were trained from scratch. Except DoWhy, the MMD distance with reconstructed samples from the methods are close to oracle performance.
Method MMD($\widehat{D^{\text{sample}}_{\bm{X}}}$, $D^{\textnormal{test}}_{\bm{X}}$) MMD($\widehat{D^{\textnormal{train}}_{\bm{X}}}$, $D^{\textnormal{test}}_{\bm{X}}$) MMD($D^{\textnormal{train}}_{\bm{X}}$, $D^{\textnormal{test}}_{\bm{X}}$)
DoWhy $0.015$ $0.014$ $0.005$
DECI $0.014$ $0.005$ $0.005$
FiP $0.015$ $0.005$ $0.005$
Cond-FiP $0.013$ $0.005$ $0.005$
Table 4: Results for Sachs dataset. We compare Cond-FiP against the baselines for the task of generating sample data on the real world benchmark. Each cell reports the MMD, and we also report the reconstruction error for all of the methods.

### v3

@@ E.1 Experiments with $n_{\mathcal{D_{\mathrm{test}}}}=100$
We report the results for the task of noise prediction, sample generation, and interventional generation in Table 10Table 11, and Table 12 respectively. We find that Cond-FiP exhibits superior generalization as compared to baselines. For example, in the case of RFF IN, Cond-FiP is even better than FiP for all the tasks! This can be attributed to the advantage of amortized inference; as the sample size in test dataset decreases, the generalization of baselines would be affected a lot since they require training from scratch on these datasets. However, amortized inference methods would be impacted less as they do not have to trained from scratch, and the inductive bias learned by them can help them generalize even with smaller input context.
Method Total Nodes LIN IN RFF IN LIN OUT RFF OUT
DoWhy 10 $0.06$ $(0.01)$ $0.22$ $(0.03)$ $0.09$ $(0.01)$ $0.16$ $(0.03)$
DECI 10 $0.15$ $(0.01)$ $0.3$ $(0.02)$ $0.22$ $(0.01)$ $0.3$ $(0.03)$
FiP 10 $0.07$ $(0.01)$ $0.18$ $(0.01)$ $0.12$ $(0.01)$ $0.11$ $(0.01)$
Cond-FiP 10 $0.07$ $(0.01)$ $0.14$ $(0.01)$ $0.09$ $(0.01)$ $0.14$ $(0.01)$
DoWhy 20 $0.06$ $(0.01)$ $0.27$ $(0.05)$ $0.07$ $(0.01)$ $0.37$ $(0.01)$
DECI 20 $0.15$ $(0.02)$ $0.33$ $(0.02)$ $0.17$ $(0.02)$ $0.35$ $(0.03)$
FiP 20 $0.09$ $(0.01)$ $0.21$ $(0.03)$ $0.1$ $(0.01)$ $0.27$ $(0.03)$
Cond-FiP 20 $0.08$ $(0.01)$ $0.12$ $(0.01)$ $0.1$ $(0.01)$ $0.15$ $(0.01)$
DoWhy 50 $0.06$ $(0.01)$ $0.29$ $(0.04)$ $0.05$ $(0.01)$ $0.47$ $(0.04)$
DECI 50 $0.14$ $(0.01)$ $0.33$ $(0.02)$ $0.14$ $(0.02)$ $0.4$ $(0.03)$
FiP 50 $0.08$ $(0.01)$ $0.23$ $(0.03)$ $0.08$ $(0.01)$ $0.37$ $(0.04)$
Cond-FiP 50 $0.08$ $(0.0)$ $0.12$ $(0.01)$ $0.08$ $(0.01)$ $0.15$ $(0.01)$
DoWhy 100 $0.06$ $(0.01)$ $0.31$ $(0.04)$ $0.06$ $(0.01)$ $0.5$ $(0.03)$
DECI 100 $0.13$ $(0.01)$ $0.36$ $(0.03)$ $0.12$ $(0.02)$ $0.44$ $(0.02)$
FiP 100 $0.08$ $(0.01)$ $0.25$ $(0.04)$ $0.1$ $(0.01)$ $0.39$ $(0.03)$
Cond-FiP 100 $0.07$ $(0.0)$ $0.13$ $(0.01)$ $0.08$ $(0.01)$ $0.17$ $(0.01)$
Table 10: Results for Noise Prediction with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=100$). We compare Cond-FiP against the baselines for the task of predicting noise variable from input observations. Each test dataset contains $100$ samples, as opposed to $400$ samples in Table 1. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
Method Total Nodes LIN IN RFF IN LIN OUT RFF OUT
DoWhy 10 $0.1$ $(0.01)$ $0.3$ $(0.06)$ $0.12$ $(0.02)$ $0.19$ $(0.03)$
DECI 10 $0.23$ $(0.01)$ $0.45$ $(0.04)$ $0.31$ $(0.02)$ $0.38$ $(0.04)$
FiP 10 $0.13$ $(0.01)$ $0.29$ $(0.04)$ $0.18$ $(0.02)$ $0.15$ $(0.03)$
Cond-FiP 10 $0.09$ $(0.01)$ $0.2$ $(0.03)$ $0.09$ $(0.02)$ $0.14$ $(0.02)$
DoWhy 20 $0.11$ $(0.01)$ $0.47$ $(0.15)$ $0.11$ $(0.02)$ $0.5$ $(0.03)$
DECI 20 $0.26$ $(0.02)$ $0.53$ $(0.05)$ $0.26$ $(0.03)$ $0.57$ $(0.04)$
FiP 20 $0.17$ $(0.02)$ $0.34$ $(0.06)$ $0.17$ $(0.02)$ $0.39$ $(0.03)$
Cond-FiP 20 $0.08$ $(0.0)$ $0.31$ $(0.06)$ $0.13$ $(0.01)$ $0.37$ $(0.02)$
DoWhy 50 $0.11$ $(0.01)$ $0.42$ $(0.08)$ $0.09$ $(0.01)$ $0.66$ $(0.06)$
DECI 50 $0.23$ $(0.02)$ $0.59$ $(0.08)$ $0.27$ $(0.04)$ $0.73$ $(0.06)$
FiP 50 $0.13$ $(0.01)$ $0.38$ $(0.07)$ $0.14$ $(0.01)$ $0.58$ $(0.06)$
Cond-FiP 50 $0.1$ $(0.01)$ $0.32$ $(0.05)$ $0.12$ $(0.01)$ $0.54$ $(0.05)$
DoWhy 100 $0.11$ $(0.01)$ $0.44$ $(0.08)$ $0.11$ $(0.01)$ $0.74$ $(0.05)$
DECI 100 $0.25$ $(0.02)$ $0.62$ $(0.08)$ $0.25$ $(0.01)$ $0.78$ $(0.07)$
FiP 100 $0.15$ $(0.01)$ $0.4$ $(0.07)$ $0.19$ $(0.02)$ $0.67$ $(0.07)$
Cond-FiP 100 $0.11$ $(0.01)$ $0.35$ $(0.07)$ $0.14$ $(0.02)$ $0.63$ $(0.07)$
Table 11: Results for Sample Generation with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=100$). We compare Cond-FiP against the baselines for the task of generating samples from the input noise variable. Each test dataset contains $100$ samples, as opposed to $400$ samples in Table 2. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
Method Total Nodes LIN IN RFF IN LIN OUT RFF OUT
DoWhy 10 $0.09$ $(0.01)$ $0.34$ $(0.08)$ $0.11$ $(0.01)$ $0.2$ $(0.04)$
DECI 10 $0.24$ $(0.02)$ $0.43$ $(0.04)$ $0.26$ $(0.03)$ $0.35$ $(0.04)$
FiP 10 $0.13$ $(0.01)$ $0.29$ $(0.04)$ $0.14$ $(0.02)$ $0.14$ $(0.03)$
Cond-FiP 10 $0.09$ $(0.02)$ $0.21$ $(0.03)$ $0.09$ $(0.01)$ $0.12$ $(0.02)$
DoWhy 20 $0.1$ $(0.01)$ $0.37$ $(0.08)$ $0.11$ $(0.02)$ $0.49$ $(0.04)$
DECI 20 $0.25$ $(0.03)$ $0.5$ $(0.05)$ $0.28$ $(0.03)$ $0.54$ $(0.04)$
FiP 20 $0.16$ $(0.01)$ $0.33$ $(0.06)$ $0.2$ $(0.03)$ $0.38$ $(0.03)$
Cond-FiP 20 $0.1$ $(0.01)$ $0.27$ $(0.05)$ $0.15$ $(0.02)$ $0.29$ $(0.03)$
DoWhy 50 $0.12$ $(0.02)$ $0.49$ $(0.14)$ $0.09$ $(0.01)$ $0.64$ $(0.07)$
DECI 50 $0.26$ $(0.03)$ $0.56$ $(0.07)$ $0.26$ $(0.03)$ $0.72$ $(0.06)$
FiP 50 $0.16$ $(0.02)$ $0.36$ $(0.06)$ $0.15$ $(0.01)$ $0.57$ $(0.06)$
Cond-FiP 50 $0.13$ $(0.02)$ $0.29$ $(0.04)$ $0.12$ $(0.01)$ $0.49$ $(0.07)$
DoWhy 100 $0.11$ $(0.01)$ $0.46$ $(0.07)$ $0.11$ $(0.01)$ $1.16$ $(0.38)$
DECI 100 $0.24$ $(0.02)$ $0.62$ $(0.08)$ $0.26$ $(0.01)$ $0.78$ $(0.07)$
FiP 100 $0.16$ $(0.02)$ $0.39$ $(0.07)$ $0.2$ $(0.02)$ $0.66$ $(0.07)$
Cond-FiP 100 $0.12$ $(0.02)$ $0.32$ $(0.07)$ $0.13$ $(0.01)$ $0.58$ $(0.07)$
Table 12: Results for Interventional Generation with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=100$). We compare Cond-FiP against the baselines for the task of generating interventional data from the input noise variable. Each test dataset contains $100$ samples, as opposed to $400$ samples in Table 3. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
E.2 Experiments with $n_{\mathcal{D_{\mathrm{test}}}}=50$
We conduct more experiments for the smaller sample size scenarios, where decrease the sample size even further to $n_{\mathcal{D_{\mathrm{test}}}}=50$ samples. We report the results for the task of noise prediction, sample generation, and interventional generation in Table 13Table 14, and Table 15 respectively. We find that baselines perform much worse than Cond-FiP for the all different SCM distributions, highlighting the efficacy of Cond-FiP for inferring causal mechanisms when the input context has smaller sample size. Note that there were issues with training DoWhy for such a small dataset, hence we do not consider them for this scenario.
Method Total Nodes LIN IN RFF IN LIN OUT RFF OUT
DECI 10 $0.19$ $(0.02)$ $0.41$ $(0.03)$ $0.2$ $(0.02)$ $0.42$ $(0.04)$
FiP 10 $0.13$ $(0.03)$ $0.27$ $(0.03)$ $0.15$ $(0.02)$ $0.21$ $(0.03)$
Cond-FiP 10 $0.09$ $(0.01)$ $0.17$ $(0.01)$ $0.11$ $(0.01)$ $0.16$ $(0.01)$
DECI 20 $0.2$ $(0.01)$ $0.42$ $(0.03)$ $0.25$ $(0.04)$ $0.45$ $(0.05)$
FiP 20 $0.12$ $(0.01)$ $0.33$ $(0.04)$ $0.15$ $(0.02)$ $0.35$ $(0.04)$
Cond-FiP 20 $0.1$ $(0.01)$ $0.16$ $(0.01)$ $0.11$ $(0.01)$ $0.17$ $(0.01)$
DECI 50 $0.2$ $(0.02)$ $0.43$ $(0.02)$ $0.2$ $(0.03)$ $0.5$ $(0.05)$
FiP 50 $0.13$ $(0.01)$ $0.32$ $(0.03)$ $0.13$ $(0.01)$ $0.49$ $(0.05)$
Cond-FiP 50 $0.1$ $(0.01)$ $0.16$ $(0.0)$ $0.1$ $(0.01)$ $0.17$ $(0.01)$
DECI 100 $0.19$ $(0.02)$ $0.43$ $(0.03)$ $0.21$ $(0.01)$ $0.53$ $(0.02)$
FiP 100 $0.11$ $(0.01)$ $0.32$ $(0.04)$ $0.13$ $(0.01)$ $0.48$ $(0.02)$
Cond-FiP 100 $0.09$ $(0.01)$ $0.16$ $(0.01)$ $0.09$ $(0.01)$ $0.18$ $(0.01)$
Table 13: Results for Noise Prediction with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=50$). We compare Cond-FiP against the baselines for the task of predicting noise variable from input observations. Each test dataset contains $50$ samples, as opposed to $400$ samples in Table 1. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows denote the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
Method Total Nodes LIN IN RFF IN LIN OUT RFF OUT
DECI 10 $0.31$ $(0.02)$ $0.58$ $(0.05)$ $0.27$ $(0.04)$ $0.49$ $(0.07)$
FiP 10 $0.2$ $(0.03)$ $0.4$ $(0.05)$ $0.21$ $(0.03)$ $0.25$ $(0.04)$
Cond-FiP 10 $0.12$ $(0.02)$ $0.28$ $(0.03)$ $0.12$ $(0.01)$ $0.18$ $(0.03)$
DECI 20 $0.34$ $(0.02)$ $0.66$ $(0.08)$ $0.39$ $(0.07)$ $0.68$ $(0.05)$
FiP 20 $0.2$ $(0.01)$ $0.51$ $(0.08)$ $0.25$ $(0.04)$ $0.51$ $(0.02)$
Cond-FiP 20 $0.13$ $(0.01)$ $0.4$ $(0.06)$ $0.19$ $(0.02)$ $0.43$ $(0.02)$
DECI 50 $0.32$ $(0.02)$ $0.66$ $(0.06)$ $0.36$ $(0.02)$ $0.8$ $(0.06)$
FiP 50 $0.2$ $(0.01)$ $0.48$ $(0.07)$ $0.22$ $(0.02)$ $0.69$ $(0.06)$
Cond-FiP 50 $0.15$ $(0.02)$ $0.4$ $(0.05)$ $0.16$ $(0.01)$ $0.59$ $(0.06)$
DECI 100 $0.36$ $(0.04)$ $0.68$ $(0.08)$ $0.39$ $(0.03)$ $0.84$ $(0.06)$
FiP 100 $0.2$ $(0.02)$ $0.49$ $(0.09)$ $0.28$ $(0.03)$ $0.73$ $(0.07)$
Cond-FiP 100 $0.16$ $(0.01)$ $0.42$ $(0.07)$ $0.22$ $(0.01)$ $0.65$ $(0.06)$
Table 14: Results for Sample Generation with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=50$). We compare Cond-FiP against the baselines for the task of generating samples from the input noise variable. Each test dataset contains $50$ samples, as opposed to $400$ samples in Table 2. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows denote the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.
Method Total Nodes LIN IN RFF IN LIN OUT RFF OUT
DECI 10 $0.3$ $(0.03)$ $0.53$ $(0.05)$ $0.26$ $(0.04)$ $0.42$ $(0.05)$
FiP 10 $0.21$ $(0.04)$ $0.35$ $(0.04)$ $0.2$ $(0.03)$ $0.22$ $(0.03)$
Cond-FiP 10 $0.12$ $(0.01)$ $0.19$ $(0.03)$ $0.07$ $(0.01)$ $0.14$ $(0.02)$
DECI 20 $0.33$ $(0.02)$ $0.6$ $(0.06)$ $0.43$ $(0.07)$ $0.63$ $(0.04)$
FiP 20 $0.21$ $(0.02)$ $0.46$ $(0.07)$ $0.29$ $(0.04)$ $0.49$ $(0.02)$
Cond-FiP 20 $0.11$ $(0.01)$ $0.29$ $(0.06)$ $0.15$ $(0.02)$ $0.32$ $(0.03)$
DECI 50 $0.34$ $(0.02)$ $0.66$ $(0.07)$ $0.34$ $(0.02)$ $0.78$ $(0.06)$
FiP 50 $0.21$ $(0.02)$ $0.46$ $(0.07)$ $0.23$ $(0.02)$ $0.68$ $(0.06)$
Cond-FiP 50 $0.13$ $(0.02)$ $0.31$ $(0.05)$ $0.12$ $(0.02)$ $0.51$ $(0.07)$
DECI 100 $0.37$ $(0.04)$ $0.67$ $(0.08)$ $0.4$ $(0.04)$ $0.84$ $(0.06)$
FiP 100 $0.21$ $(0.02)$ $0.49$ $(0.08)$ $0.28$ $(0.03)$ $0.73$ $(0.07)$
Cond-FiP 100 $0.12$ $(0.01)$ $0.33$ $(0.07)$ $0.14$ $(0.01)$ $0.58$ $(0.07)$
Table 15: Results for Interventional Generation with Smaller Sample Size ($n_{\mathcal{D_{\mathrm{test}}}}=50$). We compare Cond-FiP against the baselines for the task of generating interventional data from the input noise variable. Each test dataset contains $50$ samples, as opposed to $400$ samples in Table 3. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes much better than the baselines in this low-data regime.